Write the equation of the line that passes through point (-2,3) with a slope of -4.
Answer by venugopalramana(3286)

Write the equation of the line that passes through point A..(X1,Y1).. (-2,3) with a slope of ...M=-4.
EQN IS
Y-Y1=M(X-X1)
Y-3=-4(X+2)
Y+4X=3-8=-5
Y+4X+5=0